OpenStudy (anonymous):

Given 13.6gcm^-3 Convert it to kgm^-3 @ganeshie8

OpenStudy (anonymous):

A 13.6x10^-3 kgm^-3 B 13.6x10^-6 kgm^-3 C 13.6x10^3 kgm^-3 D 13.6x10^6 kgm^-3

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\dfrac{gm}{cm^3}=\hspace{.33em}\\~\\ \text{use the data} \hspace{.33em}\\~\\ 1~kg=10^3~g \hspace{.33em}\\~\\ 1~g=10^{-3}~kg \hspace{.33em}\\~\\ 1~m=10^{2}~cm \hspace{.33em}\\~\\ 1~cm=10^{-2}~m \hspace{.33em}\\~\\ 1~cm^3=10^{-6}~m^3 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

\[=\frac{ 13.6\times10^{-3}kg }{ 10^{-6}m^{3} }\]

OpenStudy (anonymous):

\[=13.6\times10^{-3+6}kgm^{-3}\]

OpenStudy (anonymous):

\[=13.6\times10^3kgm^{-3}\]

OpenStudy (anonymous):

Thnx @mathmath333